OpenStudy (anonymous):

Prove ab(a+b) <= a^3 + b^3 if a,b >= 0 * <= less than or equal to * >= greater than or equal to thanks

OpenStudy (anonymous):

a^3 +b^3=(a+b)(a^2-ab+b^2) then:

OpenStudy (nikvist):

\[a^2+b^2\geq2\sqrt{a^2b^2}=2ab\] \[a^2-ab+b^2\geq ab\] \[a^2-ab+b^2\geq ab/\cdot(a+b)\] \[(a^2-ab+b^2)\cdot(a+b)\geq ab\cdot(a+b)\] \[a^3+b^3\geq ab(a+b)\]

OpenStudy (anonymous):

ab(a+b)<=(a+b)(a^2-ab+b^2)=0 ab<=a^2-ab+b^2 (a-b)^2>=0 true for any a& b